OpenStudy (anonymous):

Is there anyone that can get this? Factor the expression completely. Begin by factoring out the lowest power of each common factor. 12x−1/2 + 7x^1/2 + x^3/2

OpenStudy (anonymous):

(7x+x^1/2+12)(x^-1/2)

OpenStudy (anonymous):

\[12x ^{-1/2}+7x ^{1/2}+x ^{3/2}= x ^{-1/2}(x ^{2}+7x+12)\] = (x + 4)(x + 3)/sqrt(x).